The sum of two numbers is 15. Three times one of the numbers is 19 less than five times the other. Show your solution
Answer by stanbon(75887)

The sum of two numbers is 15. Three times one of the numbers is 19 less than five times the other.
-----------------
Equations:
x + y = 15
3x = 5y-19
----
3(15-y) = 5y - 19
45 - 3y = 5y -19
8y = 64
y = 8
------
Solve for "x":
x + 8 = 15
x = 7
============
Cheers,
Stan H.
============